OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y=9/sqrt5x+5

OpenStudy (anonymous):

is it \[y=\frac{9}{\sqrt{5x+5}}\]?

OpenStudy (anonymous):

if so you have lots of choices you could say \(g(x)=\sqrt{5x+5}\)and \(f(x)=\frac{9}{x}\) or you could say \(g(x)=5x+5\) and \(f(x)=\frac{9}{\sqrt{x}}\)

OpenStudy (anonymous):

there are lots of other possibilities too, see if you can find one

OpenStudy (anonymous):

yes thank you!